OpenStudy (anonymous):

Simplify x^(3/2)x^-2 and eliminate any negative exponents. Assume all letters denote positive numbers. A) x^3 B)1/x^3 C)1/√x D) √x answer is C please explain how you get to the answer, thank you

OpenStudy (jdoe0001):

\(\Large \bf { a^{\color{brown}{ n}}\times a^{\color{blue}{ m}}\implies a^{{\color{brown}{ n}}+{\color{blue}{ m}}}\qquad thus \\ \quad \\ \quad \\ x^{\frac{3}{2}}x^{-2}\implies x^{\frac{3}{2}-2} }\)

OpenStudy (jdoe0001):

so just add those, or subtract, those fractions, and see what you get

OpenStudy (anonymous):

oh! I get it now thank you!

OpenStudy (jdoe0001):

yw